OpenStudy (anonymous):

19x3 + (14x + 4x3) =

OpenStudy (anonymous):

\[\LARGE 19x^3+(14x+4x^3)=\] ok.. so what do you have to do , simplify? :)

OpenStudy (anonymous):

please, if you don't collaborate we can't help. :(

OpenStudy (anonymous):

Do you know english? ... Ok, you might be asked to simplify it a bit. \[\LARGE 19x^3+(14x+4x^3)=\] \[\LARGE 19x^3+14x+4x^3=\] \[\LARGE (19+4)x^3+14x=\] \[\LARGE 23x^3+14x\] or maybe... \[\LARGE 19x^3+(14x+4x^3)= (19x^3+14x)+4x^3\] :)